# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ B.3 LLM pre-training
Appendix C Proof of Theorem 4.1
C.1 Preliminaries
C.1.1 Connections to kernel regression
In this section we aim to explain the assumptions made in section 2.2, namely Assumption 2.2 and 2.3.
Our setting is highly related to the classical kernel regression. In kernel regression, we aim to learn a function within a reproducing kernel Hilbert space (RKHS) from finite data samples. We begin by reviewing the foundational setup of kernel regression and subsequently show the connections between our assumptions and the classical ones.
Let $\mathcal{X}$ be a compact topological space. Let $\rho$ be a probability measure on $\mathcal{X}\times\mathbb{R}$. We denote by $\rho_{x}$ the marginal distribution on $\mathcal{X}$. Let $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a continuous PSD kernel such that $\mathbb{E}_{x\sim\rho_{x}}[K(x,x)]<\infty$. and we denote the corresponding RKHS by $\mathcal{H}_{K}$. For some $x\in\mathcal{X}$, define $K_{x}:=K(x,\cdot)\in\mathcal{H}_{K}$.
For simplicity, we assume that there exists a function $f^{*}:\mathcal{X}\to\mathbb{R}$ such that the distribution $\rho$ of $(x,y)$ is given by
$y=f^{*}(x)+\epsilon,\quad x\sim\rho_{x},$
where $\epsilon$ is a random variable independent of $x$ and satisfies $\mathbb{E}[\epsilon]=0$, $\mathbb{E}[\epsilon^{2}]=\sigma^{2}$. We call $f^{*}$ the target function.
For any $f\in\mathcal{H}_{K}$, we define the excess risk as
$\mathcal{E}(f)=\frac{1}{2}\|f-f^{*}\|_{L^{2}_{\rho_{x}}(\mathcal{X})}^{2}=\frac{1}{2}\mathbb{E}_{x\sim\rho_{x}}[(f(x)-f^{*}(x))^{2}]$
Our objective is to minimize the excess risk. The dataset is i.i.d. drawn from the distribution $\rho$.
Given a PSD kernel $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, by Mercer’s theorem there exists a nonnegative sequence $\{\lambda_{j}\}_{j=1}^{\infty}$ and an orthonormal basis $\{e_{j}\}_{j=1}^{\infty}$ of $L^{2}_{\rho}(\mathcal{X})$ such that
$K(x,x^{\prime})=\sum_{j=1}^{\infty}\lambda_{j}e_{j}(x)e_{j}(x^{\prime}),$
where the series converges absolutely and uniformly.
Define the feature functions $\phi_{j}(x)=\lambda_{j}^{\frac{1}{2}}e_{j}(x)$, then the RHKS $\mathcal{H}_{K}$ associated with $K$ can be written as
$K(x,x^{\prime})=\sum_{j=1}^{\infty}\phi_{j}(x)\phi_{j}(x^{\prime}),\quad\mathcal{H}_{K}=\left\{\sum_{j=1}^{\infty}a_{j}\phi_{j}:\sum_{j=1}^{\infty}a_{j}^{2}<\infty\right\},$
with $\{\phi_{j}\}_{j=1}^{\infty}$ being an orthonormal basis of $\mathcal{H}_{K}$. Also we have $\mathbb{E}[\phi_{j}(x)^{2}]=\lambda_{j}\mathbb{E}[e_{j}(x)^{2}]=\lambda_{j}$, and $\{\phi_{j}\}_{j=1}^{\infty}$ is still orthogonal in $L^{2}_{\rho}(\mathcal{X})$.
The interpolation space $\mathcal{H}_{K}^{s}$ is
$\mathcal{H}_{K}^{s}=\left\{\sum_{j=1}^{\infty}a_{j}\phi_{j}:\sum_{j=1}^{\infty}\frac{a_{j}^{2}}{\lambda_{j}^{s-1}}<\infty\right\}.$
where $\mathcal{H}_{K}^{0}=L^{2}_{\rho_{x}}$, $\mathcal{H}_{K}^{1}=\mathcal{H}_{K}$. We have $\mathcal{H}_{K}^{s_{1}}\subset\mathcal{H}_{K}^{s_{2}}$ for any $s_{1}>s_{2}$.
The source-capacity condition are usually given as
Assumption C.1 (Capacity condition).
There exists some $\beta>1$ such that $\lambda_{j}=O(j^{-\beta})$.
Assumption C.2 (Source condition).
There exists some $s>0$ such that $f^{*}(x)\in\mathcal{H}_{K}^{s}$.
In Assumption 2.2 and 2.3, the parameter $\theta_{j}$ corresponds to $a_{j}$ here, therefore they are the power-law version of the standard source-capacity condition, which gives the following insights:
  • A smaller $s$ indicates that the target function $f^{*}$ lies in a larger space, hence the learning problem is more difficult.
  • A smaller $\beta$ indicates that the RKHS $\mathcal{H}_{K}$ is larger, and in our setting $\mathcal{H}_{K}$ is the hypothesis space, meaning that the model capacity is larger.
C.1.2 The SDE modeling
In our setup, the SGD update can be written as
$\mathbf{v}_{k}=\mathbf{v}_{k-1}-\varphi_{k}\nabla\mathcal{R}(\mathbf{v}_{k-1})h-\varphi_{k}h\bm{\xi}_{k}.$
The term $\bm{\xi}_{k}$ is the gradient noise, whose covariance is $\frac{1}{B_{k}}\Sigma(\mathbf{v}_{k-1})$.
Therefore in order to approximate it as an Itô-type SDE, we assume the noise is Gaussian and write
$\mathbf{v}_{k}-\mathbf{v}_{k-1}=-\varphi_{k}\nabla\mathcal{R}(\mathbf{v}_{k-1})h+\varphi_{k}\sqrt{\frac{h}{B_{k}}\Sigma(\mathbf{v}_{k-1})}\sqrt{h}\mathbf{z},\quad\mathbf{z}\sim\mathcal{N}(0,\mathbf{I}).$
Hence the corresponding SDE (5) is
where $\mathbf{B}_{\tau}\in\mathbb{R}^{M}$ denotes the $M$-dimensional Brownian motion, and $\varphi$, $b$ are the continuous version of LRS function and batch-size schedule function, respectively.
C.2 Analysis of the Intrinsic-time SDE
In this section, we will derive a Volterra integral equation of the excess risk from the intrinsic-time SDE. Recall the intrinsic-time SDE:
$\mathop{}\!\mathrm{d}\bm{\nu}_{t}=-\nabla\mathcal{R}(\bm{\nu}_{t})\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}\Sigma(\bm{\nu}_{t})}\mathop{}\!\mathrm{d}\mathbf{B}_{t},$ (14)
where $\gamma_{t}=\gamma_{\varphi,b}(t)$. Here, we drop the dependence on $\varphi$ and $b$ for simplicity.
By (2), we have $\nabla\mathcal{R}(\mathbf{v})=\mathbf{W}\mathbf{H}(\mathbf{W}^{\top}\mathbf{v}-\theta^{*}).$ Let $\mathbf{u}_{t}=\mathbf{W}^{\top}\bm{\nu}_{t}-\bm{\theta}^{*}$. Then, we have
@@ C.2 Analysis of the Intrinsic-time SDE
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}\mathbf{W}^{\top}\Sigma_{t}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{t},$ (15)
where $\Sigma_{t}:=\Sigma(\bm{\nu}_{t})$.
@@
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}$ $=\mathop{}\!\mathrm{d}(\mathbf{W}^{\top}\bm{\nu}_{t}-\mathbf{v}^{*})=\mathbf{W}^{\top}\mathop{}\!\mathrm{d}\bm{\nu}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\mathbf{W}^{\top}\sqrt{\gamma_{t}\Sigma_{t}}\mathop{}\!\mathrm{d}\tilde{\mathbf{B}}_{t}.$
@@
It is easy to see that the diffusion term $\mathbf{W}^{\top}\sqrt{\gamma_{t}\Sigma_{t}}\mathop{}\!\mathrm{d}\tilde{\mathbf{B}}_{t}$ has the same distribution as $\sqrt{\gamma_{t}\mathbf{W}^{\top}\Sigma_{t}\mathbf{W}}\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}$, hence the SDE can be written in ${\mathbf{B}}_{t}$ as
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}\mathbf{W}^{\top}\Sigma_{t}\mathbf{W}}\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}.$
@@
$(2C_{1}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\preceq\Sigma(\mathbf{v})\preceq(2C_{2}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\mathbf{W}\mathbf{H}\mathbf{W}^{\top}.$
Noting $\nabla^{2}\mathcal{R}(\mathbf{v})=\mathbf{W}\mathbf{H}\mathbf{W}^{\top}$ and $\mathcal{R}(\mathbf{v})=\mathcal{E}(\mathbf{v})+\frac{1}{2}\sigma^{2}$, this lemma means $\Sigma(\mathbf{v})\eqsim\mathcal{R}(\mathbf{v})\nabla^{2}\mathcal{R}(\mathbf{v}).$ That is, the gradient noise scales proportionally with the population risk and aligns with the local curvature. Notably, the noise has two distinct sources: (i) the fit-dependent term $\mathcal{E}(\mathbf{v})$, which arises purely from minibatching and persists even in the absence of label noise; (ii) the $\sigma^{2}$ term, which captures the contribution from label noise. This anisotropic structure of SGD noise – scaling with risk and shaped by curvature – has also been observed in prior work [68, 67].
@@
$\Sigma(\mathbf{v})$ $=\mathbb{E}[\bm{\xi}\bm{\xi}^{\top}|\mathbf{v}]$
@@
$\Sigma(\mathbf{v})\preceq C_{2}\mathbf{W}\mathrm{tr}(\mathbf{H}\mathbf{u}\mathbf{u}^{\top})\mathbf{H}\mathbf{W}^{\top}+\sigma^{2}\mathbf{W}\mathbf{H}\mathbf{W}^{\top}=\left(2C_{2}\mathcal{E}(\mathbf{u})+\sigma^{2}\right)\mathbf{W}\mathbf{H}\mathbf{W}^{\top},$
@@
where $\mathbf{A}_{t}:=e^{-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}t}$, $\mathbf{S}:=\sqrt{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\mathbf{W}}$, and $2C_{1}\leqslant c_{\tau}\leqslant 2C_{2}$ are some constants.
@@
$=e^{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}t}\sqrt{\gamma_{t}\mathbf{W}^{\top}\Sigma_{t}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{t}.$
@@
$e^{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}t}\mathbf{u}_{t}-\mathbf{u}_{0}=\int_{0}^{t}e^{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\tau}\sqrt{\gamma_{\tau}\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{\tau}.$
@@
$\mathbf{u}_{t}=\mathbf{A}_{t}\mathbf{u}_{0}+\int_{0}^{t}\mathbf{A}_{t-\tau}\sqrt{\gamma_{\tau}\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{\tau}.$
@@
$=\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}+\mathbb{E}\int_{0}^{t}\gamma_{\tau}\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\right)\mathop{}\!\mathrm{d}\tau.$
@@
$\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\right)\leqslant(2\mathcal{E}_{\tau}+\sigma^{2})\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S}),$
$\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\right)\geqslant(\mathcal{E}_{\tau}+\sigma^{2})\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S}).$
Hence there exists some constant $c_{\tau}\in[2C_{1},2C_{2}]$ such that
$\mathbb{E}\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\Sigma_{\tau}\mathbf{W}}\right)=(c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S}),$
@@
C.3 The g Functions
Our analysis will frequently use following family of functions
$g_{a}(t)=g_{a}^{(M)}(t)$ $=\int_{M^{-\beta}}^{1}u^{a-1}e^{-2ut}\mathop{}\!\mathrm{d}u$ (17)
$G_{a}(t)$ $=\lim_{M\to\infty}g_{a}^{(M)}(t)=\int_{0}^{1}u^{a-1}e^{-2ut}\mathop{}\!\mathrm{d}u$
Thus, the forgetting kernel and risk functions are special cases:
$\mathcal{K}(t)=g_{2-1/\beta}(t),\quad e(t)=g_{s}(t).$ (18)

### v3

@@ B.2 LLM pre-training
Appendix C Proofs for Section 4
C.1 Volterra Integral Equation Governing the Loss Dynamics
In this section, we derive a Volterra-type integral equation that exactly characterizes the evolution of expected loss under the intrinsic-time SDE. This equation serves as the starting point for all subsequent theoretical analysis. Recall the intrinsic-time SDE:
$\mathop{}\!\mathrm{d}\bm{\nu}_{t}=-\nabla\mathcal{R}(\bm{\nu}_{t})\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}\bm{\Sigma}(\bm{\nu}_{t})}\mathop{}\!\mathrm{d}\mathbf{B}_{t},$ (32)
where we write $\gamma_{t}=\gamma(t)$ for simplicity.
By the definition of $\mathcal{R}(\mathbf{v})$, we have $\nabla\mathcal{R}(\mathbf{v})=\mathbf{W}\mathbf{H}(\mathbf{W}^{\top}\mathbf{v}-\theta^{*}).$ Let $\mathbf{u}_{t}=\mathbf{W}^{\top}\bm{\nu}_{t}-\bm{\theta}^{*}$. Then, we have
@@ C.1 Volterra Integral Equation Governing the Loss Dynamics
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}\mathbf{W}^{\top}\bm{\Sigma}_{t}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{t},$ (33)
where $\bm{\Sigma}_{t}:=\bm{\Sigma}(\bm{\nu}_{t})$.
@@
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}$ $=\mathop{}\!\mathrm{d}(\mathbf{W}^{\top}\bm{\nu}_{t}-\mathbf{v}^{*})=\mathbf{W}^{\top}\mathop{}\!\mathrm{d}\bm{\nu}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\mathbf{W}^{\top}\sqrt{\gamma_{t}\bm{\Sigma}_{t}}\mathop{}\!\mathrm{d}\tilde{\mathbf{B}}_{t}.$
@@
It is easy to see that the diffusion term $\mathbf{W}^{\top}\sqrt{\gamma_{t}\bm{\Sigma}_{t}}\mathop{}\!\mathrm{d}\tilde{\mathbf{B}}_{t}$ has the same distribution as $\sqrt{\gamma_{t}\mathbf{W}^{\top}\bm{\Sigma}_{t}\mathbf{W}}\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}$, hence the SDE can be written in ${\mathbf{B}}_{t}$ as
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}\mathbf{W}^{\top}\bm{\Sigma}_{t}\mathbf{W}}\mathop{}\!\mathrm{d}{\mathbf{B}}_{t}.$
@@
$(2\rho_{-}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\preceq\bm{\Sigma}(\mathbf{v})\preceq(2\rho_{+}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\mathbf{W}\mathbf{H}\mathbf{W}^{\top}.$
Noting $\nabla^{2}\mathcal{R}(\mathbf{v})=\mathbf{W}\mathbf{H}\mathbf{W}^{\top}$ and $\mathcal{R}(\mathbf{v})=\mathcal{E}(\mathbf{v})+\frac{1}{2}\sigma^{2}$, this lemma means $\bm{\Sigma}(\mathbf{v})\eqsim\mathcal{R}(\mathbf{v})\nabla^{2}\mathcal{R}(\mathbf{v}).$ That is, the gradient noise scales proportionally with the population risk and aligns with the local curvature. Notably, the noise has two distinct sources: (i) the fit-dependent term $\mathcal{E}(\mathbf{v})$, which arises purely from minibatching and persists even in the absence of label noise; (ii) the $\sigma^{2}$ term, which captures the contribution from label noise. This anisotropic structure of SGD noise – scaling with risk and shaped by curvature – has also been observed in prior work [69, 68].
@@
$\bm{\Sigma}(\mathbf{v})$ $=\mathbb{E}[\bm{\xi}\bm{\xi}^{\top}|\mathbf{v}]$
@@
$\bm{\Sigma}(\mathbf{v})\preceq\rho_{+}\mathbf{W}\mathrm{tr}(\mathbf{H}\mathbf{u}\mathbf{u}^{\top})\mathbf{H}\mathbf{W}^{\top}+\sigma^{2}\mathbf{W}\mathbf{H}\mathbf{W}^{\top}=\left(2\rho_{+}\mathcal{E}(\mathbf{u})+\sigma^{2}\right)\mathbf{W}\mathbf{H}\mathbf{W}^{\top},$
@@
where $\mathbf{A}_{t}:=e^{-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}t}$, $\mathbf{S}:=\sqrt{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\mathbf{W}}$, and $c_{\tau}\in[2\rho_{-},2\rho_{+}]$ for any $\tau\geqslant 0$.
@@
$=e^{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}t}\sqrt{\gamma_{t}\mathbf{W}^{\top}\bm{\Sigma}_{t}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{t}.$
@@
$e^{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}t}\mathbf{u}_{t}-\mathbf{u}_{0}=\int_{0}^{t}e^{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\tau}\sqrt{\gamma_{\tau}\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{\tau}.$
@@
$\mathbf{u}_{t}=\mathbf{A}_{t}\mathbf{u}_{0}+\int_{0}^{t}\mathbf{A}_{t-\tau}\sqrt{\gamma_{\tau}\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{\tau}.$
@@
$=\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}+\mathbb{E}\int_{0}^{t}\gamma_{\tau}\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\right)\mathop{}\!\mathrm{d}\tau.$
@@
$\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\right)\leqslant(2\rho_{+}\mathcal{E}_{\tau}+\sigma^{2})\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S}),$
$\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\right)\geqslant(2\rho_{-}\mathcal{E}_{\tau}+\sigma^{2})\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S}).$
Hence there exists some constant $c_{\tau}\in[2\rho_{-},2\rho_{+}]$ such that
$\mathbb{E}\mathrm{tr}\left(\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\sqrt{\mathbf{W}^{\top}\bm{\Sigma}_{\tau}\mathbf{W}}\right)=(c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S}),$
@@
C.2 Power-law Decay of the $e_{M}$ and $\mathcal{K}_{M}$ Functions
In this section, we rigorously establish the power-law decay behavior of the functions $e_{M}$ and $\mathcal{K}_{M}$.